# Assessing HRV and HR Dynamics with Wearables During Socially Anxious Situations: Insights from a Controlled Study in a Low-Middle-Income Country

### v1

@@ 2.4. Wearables and smartphones for mental disorder detection
(miranda2014anxiety, ) Use wearables to detect anxiety level 4 8 Presentation (10 min) HR Between group Insignificant HR increase in the mild SAD group
@@
(gaebler2013heart, ) Investigate HRV in SAD patients 22 22 Resting (5min) and Emotional face matching task (approx. 5min) HR, HF, LF/HF Between group Rest: Significantly reduced HF and increased LF/HF ratio in SAD group but no difference in HR. Task: Significantly less HF in SAD group.
@@
(tolin2021psychophysiological, ) Examine psychophysiological arousal variations in different phases 16 28 Baseline (3 min), stressor (6 min), Recovery (3 min) HF, RMSSD, HR Mixed linear model Increased HR in SAD compared to control. Reduced HF in SAD during stressor. No significant changes in RMSSD
@@
(garcia2017autonomic, ) Study how SAD individuals react before, during, and after the TSST activity 18 21 Baseline (10 min), TSST activity (5 min), Recovery (35 min) HR, RMSSD, LF/HF Both between and within Significantly higher HR and lower RMSSD in SAD individuals compared to controls.
@@
(tamura2013salivary, ) Compare HRV of SAD individuals with controls 32 80 Stimulation (40s) LF, HF, LF/HF, HR Between No significant difference in HR and HRV between SAD and non-SAD groups.
@@
(bailey2019moderating, ) Examine HRV in SAD individuals in a 24-hour study 16 16 24hr (Lying, Sitting, Standing, Walking, Running, Wake, Sleep) RMSSD, HR General linear mixed modeling No significant difference between HR and HRV during the wake. Further, they found higher HR during sleep. However no group difference in RMSSD during sleep.
@@
(harrewijn2018heart, ) Investigate whether HRV during rest and social tasks is a candidate endophenotype of SAD 17 104 Resting state (3 min), anticipation (5 min), speech (3 min), recovery (5 min), and second resting state (5 min) RMSSD, HF, LF/HF, HR Regression model Participants with SAD exhibited lower HRV compared to controls. However, the results were not significant.
@@
(madison2021social, ) Examine association between SAD symptoms and HRV 124 - Resting state (5 min), Vocal emotion recognition (2 min) RMSSD Linear regression Individuals with greater social anxiety symptoms have significantly lower task HRV but not lower resting HRV.
@@
(grossman2001gender, ) Examine cardiovascular responses in older people in social situations 30 30 Baseline (10min), paced breathing (5 min), speech preparation (4 min), speech (4 min) HR MANOVA Significantly increased HR for both the groups from baseline to speech preparation and speech presentation. Women had significantly increased HR than men in both tasks.
@@
(pittig2013heart, ) Examine HF-HRV in SAD and other anxiety disorder participants 25 39 Baseline (5 min), relaxation (15 min), hyperventilation (1 min) HR and HF Hierarchical linear modeling No significant difference in HR in baseline, relaxation, and hyperventilation between SAD and non-SAD. However, non-SAD groups show significantly higher HF in the baseline, with no difference in relaxation and hyperventilation.
@@
(held2021heart, ) Investigate HRV and HR in clinically anxious participants 26 14 Baseline (3 min), Working memory Task (10 min), Recovery (3 min) HF and HR Multilevel modeling No significant difference in HF and HRV at baseline and reflection phase.
3. Methodology
Table 2. Participants characteristics
@@ 3.2. Study procedure
Figure 2. Participants participating in the controlled lab study
Figure 3. Sequence of different phases in the study. PAL denotes instances at which participants reported perceived anxiety levels (PAL) through surveys. Shaded blocks Refer to caption, represent instances whenever RA gave instructions to the participants.
@@
  3. (3)
    Activity Phase: During the activity phase, each participant gave a speech in front of the other two participants and RA for 2.5 minutes on an assigned topic. We chose speech as our anxiety-provoking activity while following the Trier Social Stress Test (TSST) guidelines. Speech is a well-known TSST activity like email-usage pattern (akbar2019email, ) or math test (paredes2018fast, ) and has been used to investigate the changes in the HR/HRV parameters (garcia2017autonomic, ; harrewijn2018heart, ; mauss2003autonomic, ; mauss2004there, ). The RA instructions to P1 were, “Your speech topic is $<$assigned topic$>$. Take 30 seconds to prepare the speech”. After 30 seconds, the RA asks P1 “please start your Speech”. Following the speech activity, P1 reported her PAL by completing a survey. Similarly, P2 and P3 complete the speech activity and survey forms one after the other. The speech topics were, “Importance of Value Education”, “Animal Zoos should be banned”, “Effect of fake news on society”, etc.
  4. (4)
@@
During the study, only one participant refused to continue with the study after reading the PIS. For each study session, we had two dummy participants as a backup to handle such situations where either a participant refuses to take part or does not appear on time. Dummy participants participated in seven study sessions to ensure study homogeneity. The dummy participant’s data was later excluded during the data analysis. It took us four months to collect the data from a total of 99 participants.
3.3. Ground Truth
@@
3.3.1. Social Phobia Inventory (SPIN)
Social anxiety can manifest in various forms (such as physical, psychological, and behavioral) across individuals in social situations (stein2008social, ; ollendick2002developmental, ). We used the SPIN questionnaire to assess the severity of SAD in participants. It is effective in screening and measuring the severity of SAD and is widely used for participant screening worldwide (connor2000psychometric, ; miranda2014anxiety, ; chukwujekwu2018validation, ). Developed by the Department of Psychiatry and Behavioral Sciences at Duke University, SPIN consists of 17 items addressing issues such as “fear of physical symptoms”, “fear of negative evaluation”, and “fear of uncertainty in social situations”. Participants were requested to respond to all 17 questions on a Likert scale from 0 to 4, where 0 indicated “Not at all” and 4 indicated “Extremely”. The SPIN responses were collected before the placement of sensors, and participants were unaware of the activities they would later perform. Further, we used the SPIN responses during data analysis to understand the severity of the SAD of the participants and classify them into SAD and non-SAD groups. Figure 4 shows the distribution of SPIN scores at a bin width of 3.
3.3.2. Perceived Anxiety Level (PAL)
@@
Figure 4. Distribution of SPIN Scores. Participants with SPIN scores greater than 20 were labeled as SAD, while those with scores of 20 or below were labeled as non-SAD.
@@
Figure 5. Distribution of perceived anxiety level (PAL) at the baseline and different phases (anticipation, speech activity, and reflection).
3.4. Study rationale
@@
Figure 6. Figures showing Raw ECG signal, filtered ECG signal, and filtered signal with R peaks.
4. Analysis
4.1. Data cleaning
Before analyzing ECG data, filtering out noise and artifacts caused by equipment issues, body movements, or poor electrode contact is crucial. To achieve this, we applied a Finite Impulse Response (FIR) filter with a bandpass frequency range of 1-49 Hz to the recorded ECG signals. We compared the filtered signals to the ideal ECG waveform to evaluate the filter’s effectiveness and found consistent results across all participants’ data. Additionally, we utilized the Hamilton Segmenter algorithm for R peak detection (hamilton2002open, ). Figure 6 illustrates the raw, filtered, and R peak detection plots for one participant’s ECG data.
@@ 4.1. Data cleaning
4.2. Feature extraction
[…]
To extract HRV features (i.e., time, frequency, and nonlinear indices of HRV), we used the Neurokit Python package (makowski2021neurokit2, ). We fed the R-peak indices, obtained using the Hamilton Segmenter on a cleaned ECG signal, into the respective functions for time, frequency, and nonlinear indices separately. These functions provided aggregated time domain, frequency domain, and nonlinear indices for each ECG signal. This process was repeated for each participant across all phases (i.e., baseline, anticipation, activity, and reflection). As a result, we obtained the time domain, frequency domain, and nonlinear indices for each phase.
Table 3. List of HRV and HR parameters used in this study.
Features Explanation
@@ 4.2. Feature extraction
(a) Signature of ECG signal due to faulty device
(b) Signature of noisy ECG Signal
Figure 7. Examples of ECG signatures dropped during signal filtering from the study
4.3. Participant’s grouping
Following existing studies (connor2000psychometric, ; miranda2014anxiety, ; chukwujekwu2018validation, ; lauria2016primary, ), we categorized 51 participants into SAD and non-SAD groups, using a threshold of 20 on the SPIN score. Participants with scores exceeding 20 were labeled as SAD, while those with scores below were classified as non-SAD. Consequently, the total participant pool of 51 participants was divided into two distinct groups: 31 SAD and 20 non-SAD participants. Based on SPIN scores and at a significance level of 0.05, the computed effect size difference was found to be 2.68. This substantial effect size indicates a significant distinction between the two groups. The sunburst chart in Figure 8 represents the participant distribution across different levels. There were 31 SAD and 20 non-SAD participants. Among the SAD group, there were 28 males and eight females, while the non-SAD group had 11 males and nine females. Additionally, the distribution of participants’ home residence within the SAD group was as follows: 24 participants resided in urban areas and seven in rural areas, while in the non-SAD group, 19 participants were from urban areas and seven from rural areas.
Figure 8. Distribution of 51 participants.
4.4. Correlation analysis
@@
  • Correlation results of HR/HRV features of each phase with SPIN - The correlation analysis revealed that most HRV features exhibited negative correlations with the SPIN score, regardless of the phase. Specifically, MeanNN, MedianNN, Prc20NN, HF, HFn, and LnHF displayed correlations between -0.1 and -0.40, while other features fell between -0.09 and +0.09. Heart rate (HR) showed positive correlations with SPIN across all phases, with correlations ranging from +0.19 to +0.20. These findings suggest that participants with higher SPIN scores tend to have lower HRV and higher HR, highlighting a consistent relationship between social anxiety levels and these physiological measures syncing with the literature that SAD participants exhibit lower HRV and higher HR (alvares2013reduced, ).
@@
    Figure 9. Relationship between SPIN and PAL in different phases (Baseline, Anticipation, Speech, and Reflection). The X-axis represents the SPIN score, and the Y-axis represents the PAL.
@@
Figure 10. Each circle represents a cluster, and each dot represents a participant. Red, orange, green, blue, and purple dot colors at “Baseline” denote the perceived anxiety levels of 1, 2, 3, 4, and 5, respectively. The figure highlights the transition of participants’ anxiety levels from the Baseline to a specific phase (i.e., Anticipation, Speech, and Reflection). [Best viewed in color]
4.4.1. Ablation Study
This ablation study aimed to determine whether the correlation between HRV and self-reports (SPIN and PAL) differs between males and females. The rationale behind this investigation was motivated by existing literature, which suggests that females typically have lower HRV and higher HR than males (geovanini2020age, ). To explore this, we first conducted a correlation analysis between SPIN and HRV for males, followed by a similar analysis for females. We repeated a similar correlation analysis between males’ PAL and HRV followed by females.
@@ 4.4.1. Ablation Study
4.5. Hypothesis testing
The collected ECG data with demographic information (i.e., gender and location) has a multilevel structure. So, to address our hypotheses, we have used a multilevel modeling approach. Multilevel modeling, also known as hierarchical linear modeling or mixed-effects modeling, is a statistical model used to analyze data that follows a nested, clustered, or hierarchical structure (leyland2001multilevel, ; bristol_multilevel, ; leyland2020multilevel, ). It allows for estimating both within and between-group effects, capturing the complex relationships within the data. The predictor variables in such models can be modeled as fixed or random effect variables. A fixed effect variable represents the average relationship between the predictor and the response variable, assuming this relationship is consistent across all groups or levels in the data. In contrast, a random effect variable accounts for variability between groups or levels, capturing how the relationship between the predictor and the response variable might differ across these groups or levels. We performed multilevel statistical analyses using the ‘lme4.0’ package of R language (bates2014fitting, ).
@@ 4.5. Hypothesis testing
Hypothesis 1: We tested our hypothesis 1 with the following model:
@@
The results show a significant decrease in HRV features (SDNN, HTI, TINN, HF, SD1, and ApEn) and significant increases in HRV feature ‘S’ during the anticipation phase compared to the baseline. Similarly, we observed significant reductions in HRV features (MeanNN, RMSSD, MedianNN, Prc20NN, Prc80NN, pNN50, pNN20, LnHF, S, SD1, SD1SD2) and a significant increase in HR during the speech phase compared to the baseline. Conversely, during the reflection phase, we found a significant increase in HRV features (MeanNN, MedianNN, Prc20NN, HF, HFn, LnHF) and a significant decrease in HRV features (HTI, TINN, and ApEn) compared to the Baseline.
@@
Table 4. Results of hypothesis 1, 2, and 3, where (i) Acronyms ATP and RP denote Anticipation and Reflection, respectively, (ii) Arrows, $\uparrow$ and $\downarrow$ represent increase and decrease, respectively, (iii) Significance levels ‘***’, ‘**’, ‘*’, ‘.’ correspond to 0.001, 0.01, 0.05, and 0.1, respectively, (iv) $\leftrightarrow$ means that direction (increase or decrease) not reported by authors, and “-” means no literature exists.
@@
RMSSD $\downarrow$ $\downarrow$** $\downarrow$
@@
TINN $\downarrow$*** $\downarrow$ $\downarrow$***
HF $\downarrow$** $\downarrow$. $\uparrow$***
HFn $\uparrow$ $\uparrow$ $\uparrow$***
LnHF $\uparrow$ $\downarrow$** $\uparrow$***
@@
SD2 $\downarrow$. $\downarrow$ $\downarrow$
@@
SDNN $\downarrow$. $\downarrow$ (alvares2013reduced, )
RMSSD $\downarrow$* $\downarrow$ (alvares2013reduced, ; garcia2017autonomic, ; harrewijn2018heart, ; madison2021social, ), $\leftrightarrow$ (tolin2021psychophysiological, ; tamura2013salivary, ; bailey2019moderating, )
@@
HF $\uparrow$ $\downarrow$ (alvares2013reduced, ; pittig2013heart, ; gaebler2013heart, ; tolin2021psychophysiological, ; harrewijn2018heart, )
@@
HR $\uparrow$ $\uparrow$ (miranda2014anxiety, ; mauss2003autonomic, ; garcia2017autonomic, ; tolin2021psychophysiological, ; bailey2019moderating, ; grossman2001gender, ), $\leftrightarrow$ (gaebler2013heart, ; tamura2013salivary, ; pittig2013heart, )
@@
PAL $\uparrow$*** $\uparrow$*** $\downarrow$ $\uparrow$ ** $\uparrow$ (mauss2003autonomic, ; garcia2017autonomic, ; mauss2004there, )
@@
Hypothesis 2: We tested our hypothesis 2 with the following model:
@@
In this model, we treated phase and group as fixed variables, while the remaining variables were considered random effects. The baseline in phase and non-SAD in group was set as the reference. We ran the above model for each HRV feature and HR separately. Our results in Table 4(b) shows that most HRV features decreased in the SAD participants compared to non-SAD participants. However, only RMSSD and SD1 were found significant at a 0.05 significance level. Furthermore, we found a increase HR in SAD group however the result was not significant enough to be accepted.
Hypothesis 3: Hypothesis 3 focuses on PAL instead of the HRV/HR features. Specifically, it examines how PAL changes across different phases and whether SAD participants consistently report higher PAL. Here, we considered two sub-cases:

### v2

@@ 2.3 Wearables and smartphones for mental disorder detection
[miranda2014anxiety] Use wearables to detect anxiety level 4 8 Presentation (10 min) HR Between group Insignificant HR increase in the mild SAD group
@@
[gaebler2013heart] Investigate HRV in SAD patients 22 22 Resting (5min) and Emotional face matching task (approx. 5min) HR, HF, LF/HF Between group Rest: Significantly reduced HF and increased LF/HF ratio in SAD group but no difference in HR. Task: Significantly less HF in SAD group.
@@
[tolin2021psychophysiological] Examine psychophysiological arousal variations in different phases 16 28 Baseline (3 min), stressor (6 min), Recovery (3 min) HF, RMSSD, HR Mixed linear model Increased HR in SAD compared to control. Reduced HF in SAD during stressor. No significant changes in RMSSD
@@
[garcia2017autonomic] Study how SAD individuals react before, during, and after the TSST activity 18 21 Baseline (10 min), TSST activity (5 min), Recovery (35 min) HR, RMSSD, LF/HF Both between and within Significantly higher HR and lower RMSSD in SAD individuals compared to controls.
@@
[tamura2013salivary] Compare HRV of SAD individuals with controls 32 80 Stimulation (40s) LF, HF, LF/HF, HR Between No significant difference in HR and HRV between SAD and non-SAD groups.
@@
[bailey2019moderating] Examine HRV in SAD individuals in a 24-hour study 16 16 24hr (Lying, Sitting, Standing, Walking, Running, Wake, Sleep) RMSSD, HR General linear mixed modeling No significant difference between HR and HRV during the wake. Further, they found higher HR during sleep. However no group difference in RMSSD during sleep.
@@
[harrewijn2018heart] Investigate whether HRV during rest and social tasks is a candidate endophenotype of SAD 17 104 Resting state (3 min), anticipation (5 min), speech (3 min), recovery (5 min), and second resting state (5 min) RMSSD, HF, LF/HF, HR Regression model Participants with SAD exhibited lower HRV compared to controls. However, the results were not significant.
@@
[madison2021social] Examine association between SAD symptoms and HRV 124 - Resting state (5 min), Vocal emotion recognition (2 min) RMSSD Linear regression Individuals with greater social anxiety symptoms have significantly lower task HRV but not lower resting HRV.
@@
[grossman2001gender] Examine cardiovascular responses in older people in social situations 30 30 Baseline (10min), paced breathing (5 min), speech preparation (4 min), speech (4 min) HR MANOVA Significantly increased HR for both the groups from baseline to speech preparation and speech presentation. Women had significantly increased HR than men in both tasks.
@@
[pittig2013heart] Examine HF-HRV in SAD and other anxiety disorder participants 25 39 Baseline (5 min), relaxation (15 min), hyperventilation (1 min) HR and HF Hierarchical linear modeling No significant difference in HR in baseline, relaxation, and hyperventilation between SAD and non-SAD. However, non-SAD groups show significantly higher HF in the baseline, with no difference in relaxation and hyperventilation.
@@
[held2021heart] Investigate HRV and HR in clinically anxious participants 26 14 Baseline (3 min), Working memory Task (10 min), Recovery (3 min) HF and HR Multilevel modeling No significant difference in HF and HRV at baseline and reflection phase.
3 Methodology
Table 2: Participants characteristics
@@ 3.2 Study procedure
Figure 2: Participants participating in the controlled lab study
Figure 3: Sequence of different phases in the study. PAL denotes instances at which participants reported perceived anxiety levels (PAL) through surveys. Shaded blocks Refer to caption, represent instances whenever RA gave instructions to the participants.
@@
  3. 3.
    Activity Phase: During the activity phase, each participant gave a speech in front of the other two participants and RA for 2.5 minutes on an assigned topic. We chose speech as our anxiety-provoking activity while following the Trier Social Stress Test (TSST) guidelines. Speech is a well-known TSST activity like email-usage pattern [akbar2019email] or math test [paredes2018fast] and has been used to investigate the changes in the HR/HRV parameters [garcia2017autonomic, harrewijn2018heart, mauss2003autonomic, mauss2004there]. The RA instructions to P1 were, “Your speech topic is <<assigned topic>>. Take 30 seconds to prepare the speech”. After 30 seconds, the RA asks P1 “please start your Speech”. Following the speech activity, P1 reported her PAL by completing a survey. Similarly, P2 and P3 complete the speech activity and survey forms one after the other. The speech topics were, “Importance of Value Education”, “Animal Zoos should be banned”, “Effect of fake news on society”, etc.
  4. 4.
@@
During the study, only one participant refused to continue with the study after reading the PIS. For each study session, we had two dummy participants as a backup to handle such situations where either a participant refuses to take part or does not appear on time. Dummy participants participated in seven study sessions to ensure study homogeneity. The dummy participant’s data was later excluded during the data analysis. It took us four months to collect the data from a total of 99 participants. With this paper, we release our collected dataset [^9] https://osf.io/phze7/ publicly.
3.3 Ground Truth
@@
3.3.1 Social Phobia Inventory (SPIN)
Social anxiety can manifest in various forms (such as physical, psychological, and behavioral) across individuals in social situations [stein2008social, ollendick2002developmental]. We used the SPIN questionnaire to assess the severity of SAD in participants. It is effective in screening and measuring the severity of SAD and is widely used for participant screening worldwide [connor2000psychometric, miranda2014anxiety, chukwujekwu2018validation]. Developed by the Department of Psychiatry and Behavioral Sciences at Duke University, SPIN consists of 17 items addressing issues such as “fear of physical symptoms”, “fear of negative evaluation”, and “fear of uncertainty in social situations”. Participants were requested to respond to all 17 questions on a Likert scale from 0 to 4, where 0 indicated “Not at all” and 4 indicated “Extremely”. The SPIN responses were collected before the placement of sensors, and participants were unaware of the activities they would later perform. Further, we used the SPIN responses during data analysis to understand the severity of the SAD of the participants and classify them into SAD and non-SAD groups. Figure 4 shows the distribution of SPIN scores at a bin width of 4.
3.3.2 Perceived Anxiety Level (PAL)
Figure 4: Distribution of SPIN Scores. Participants with SPIN scores greater than 20 were labeled as SAD, while those with scores of 20 or below were labeled as non-SAD.
@@ 3.3.2 Perceived Anxiety Level (PAL)
Figure 5: Distribution of perceived anxiety level (PAL) at the baseline and different phases (anticipation, speech activity, and reflection).
3.4 Study rationale
@@
Figure 6: Figures showing Raw ECG signal, filtered ECG signal, and filtered signal with R peaks.
(a) Signature of ECG signal due to faulty device
(b) Signature of noisy ECG Signal
Figure 7: Examples of ECG signatures dropped during signal filtering from the study
4 Analysis
4.1 Data cleaning
Before analyzing ECG data, filtering out noise and artifacts caused by equipment issues, body movements, or poor electrode contact is crucial. To achieve this, we applied a Finite Impulse Response (FIR) filter with a bandpass frequency range of 1-49 Hz to the recorded ECG signals. We compared the filtered signals to the ideal ECG waveform to evaluate the filter’s effectiveness and found consistent results across all participants’ data. Additionally, we utilized the Hamilton Segmenter algorithm for R peak detection [hamilton2002open]. Figure 6 illustrates the raw, filtered, and R peak detection plots for one participant’s ECG data.
@@ 4.1 Data cleaning
Table 3: List of HRV and HR parameters used in this study.
Feature Explanation
@@
4.2 Feature extraction
[…]
To extract HRV features (i.e., time, frequency, and nonlinear indices of HRV), we used the Neurokit Python package [makowski2021neurokit2]. We fed the R-peak indices, obtained using the Hamilton Segmenter on a cleaned ECG signal, into the respective functions for time, frequency, and nonlinear indices separately. These functions provided aggregated time domain, frequency domain, and nonlinear indices for each ECG signal. This process was repeated for each participant across all phases (i.e., baseline, anticipation, activity, and reflection). As a result, we obtained the time domain, frequency domain, and nonlinear indices for each phase.
4.3 Participant’s grouping
Following existing studies [connor2000psychometric, miranda2014anxiety, chukwujekwu2018validation, lauria2016primary], we categorized 51 participants into SAD and non-SAD groups, using a threshold of 20 on the SPIN score. Participants with scores exceeding 20 were labeled as SAD, while those with scores below were classified as non-SAD. Consequently, the total participant pool of 51 participants was divided into two distinct groups: 31 SAD and 20 non-SAD participants. Based on SPIN scores and at a significance level of 0.05, the computed effect size difference was found to be 2.68. This substantial effect size indicates a significant distinction between the two groups.
Figure 8: Distribution of 51 participants.
The sunburst chart in Figure 8 represents the participant distribution across different levels. There were 31 SAD and 20 non-SAD participants. Among the SAD group, there were 28 males and eight females, while the non-SAD group had 11 males and nine females. Additionally, the distribution of participants’ home residence within the SAD group was as follows: 24 participants resided in urban areas and seven in rural areas, while in the non-SAD group, 19 participants were from urban areas and seven from rural areas.
4.4 Correlation analysis
@@
  • Correlation results of HR/HRV features of each phase with SPIN - The correlation analysis revealed that most HRV features exhibited negative correlations with the SPIN score, regardless of the phase. Specifically, MeanNN, MedianNN, Prc20NN, HF, HFn, and LnHF displayed correlations between -0.1 and -0.40, while other features fell between -0.09 and +0.09. Heart rate (HR) showed positive correlations with SPIN across all phases, with correlations ranging from +0.19 to +0.20. These findings suggest that participants with higher SPIN scores tend to have lower HRV and higher HR, highlighting a consistent relationship between social anxiety levels and these physiological measures syncing with the literature that SAD participants exhibit lower HRV and higher HR [alvares2013reduced].
@@
    Figure 9: Relationship between SPIN and PAL in different phases (Baseline, Anticipation, Speech, and Reflection). The X-axis represents the SPIN score, and the Y-axis represents the PAL.
@@
Figure 10: Each circle represents a cluster, and each dot represents a participant. Red, orange, green, blue, and purple dot colors at “Baseline” denote the perceived anxiety levels of 1, 2, 3, 4, and 5, respectively. The figure highlights the transition of participants’ anxiety levels from the Baseline to a specific phase (i.e., Anticipation, Speech, and Reflection). [Best viewed in color]
4.4.1 Ablation Study
This ablation study aimed to determine whether the correlation between HRV and self-reports (SPIN and PAL) differs between males and females. The rationale behind this investigation was motivated by existing literature, which suggests that females typically have lower HRV and higher HR than males [geovanini2020age]. To explore this, we first conducted a correlation analysis between SPIN and HRV for males, followed by a similar analysis for females. We repeated a similar correlation analysis between males’ PAL and HRV followed by females.
@@ 4.4.1 Ablation Study
4.5 Hypothesis testing
The collected ECG data with demographic information (i.e., gender and location) has a multilevel structure. So, to address our hypotheses, we have used a multilevel modeling approach. Multilevel modeling, also known as hierarchical linear modeling or mixed-effects modeling, is a statistical model used to analyze data that follows a nested, clustered, or hierarchical structure [leyland2001multilevel, bristol_multilevel, leyland2020multilevel]. It allows for estimating both within and between-group effects, capturing the complex relationships within the data. The predictor variables in such models can be modeled as fixed or random effect variables. A fixed effect variable represents the average relationship between the predictor and the response variable, assuming this relationship is consistent across all groups or levels in the data. In contrast, a random effect variable accounts for variability between groups or levels, capturing how the relationship between the predictor and the response variable might differ across these groups or levels. We performed multilevel statistical analyses using the ‘lme4.0’ package of R language [bates2014fitting].
@@ 4.5 Hypothesis testing
H1: We tested our hypothesis 1 with the following model:
@@
The results show a significant decrease in HRV features (SDNN, HTI, TINN, HF, SD1, and ApEn) and significant increases in HRV features HF, HFn, and LnHF during the anticipation phase compared to the baseline. Similarly, we observed significant reductions in HRV features (MeanNN, RMSSD, MedianNN, Prc20NN, pNN20, TINN, LnHF, S, SD1, and SD1SD2) and a significant increase in HR during the speech phase compared to the baseline. Conversely, during the reflection phase, we found a significant increase in HRV features (MeanNN, MedianNN, Prc20NN, HF, HFn, LnHF) and a significant decrease in HRV features (HTI, TINN, and ApEn) compared to the Baseline.
@@
Table 4: Results of hypothesis 1, 2, and 3, where (i) Acronyms ATP and RP denote Anticipation and Reflection, respectively, (ii) Arrows, $\uparrow$ and $\downarrow$ represent increase and decrease, respectively, (iii) Significance levels ‘***’, ‘**’, ‘*’, ‘.’ correspond to 0.001, 0.01, 0.05, and 0.1, respectively, (iv) $\leftrightarrow$ means that direction (increase or decrease) not reported by authors, and “-” means no literature exists.
@@
RMSSD $\downarrow$. $\downarrow**$ $\downarrow$
@@
TINN $\downarrow***$ $\downarrow$. $\downarrow***$
HF $\uparrow**$ $\downarrow$. $\uparrow***$
HFn $\uparrow***$ $\uparrow$ $\uparrow***$
LnHF $\uparrow*$ $\downarrow**$ $\uparrow**$
@@
SD2 $\downarrow**$ $\downarrow$ $\downarrow$
@@
SDNN $\downarrow$. $\downarrow$ [alvares2013reduced]
RMSSD $\downarrow*$ $\downarrow$ [alvares2013reduced, garcia2017autonomic, harrewijn2018heart, madison2021social], $\leftrightarrow$ [tolin2021psychophysiological, tamura2013salivary, bailey2019moderating]
@@
HF $\downarrow$ $\downarrow$ [alvares2013reduced, pittig2013heart, gaebler2013heart, tolin2021psychophysiological, harrewijn2018heart]
@@
HR $\uparrow$ $\uparrow$ [miranda2014anxiety, mauss2003autonomic, garcia2017autonomic, tolin2021psychophysiological, bailey2019moderating, grossman2001gender], $\leftrightarrow$ [gaebler2013heart, tamura2013salivary, pittig2013heart]
@@
PAL $\uparrow***$ $\uparrow***$ $\downarrow$ $\uparrow$ ** $\uparrow$ [mauss2003autonomic, garcia2017autonomic, mauss2004there]
@@
H2: We tested our hypothesis 2 with the following model:
@@
In this model, we treated phase and group as fixed variables, while the remaining variables were considered random effects. The baseline in phase and non-SAD in group was set as the reference. We ran the above model for each HRV feature and HR separately. Our result in Table 4(b) shows that most HRV features decreased in the SAD participants compared to non-SAD participants irrespective of the phases. However, only RMSSD, S, and SD1 were found significant at a 0.05 significance level. Furthermore, we found an increase in HR in the SAD group; however, the result was not significant enough to be accepted.
H3: Hypothesis 3 focuses on PAL instead of the HRV/HR features. Specifically, it examines how PAL changes across different phases and whether SAD participants consistently report higher PAL. Here, we considered two sub-cases: